# Predictive Communications for Low-Altitude Networks

### v1

@@ Abstract
The emergence of dense, mission-driven aerial networks supporting the low-altitude economy presents unique communication challenges, including extreme channel dynamics and severe cross-tier interference. Traditional reactive communication paradigms are ill-suited to these environments, as they fail to leverage the network’s inherent predictability. This paper introduces predictive communication, a novel paradigm transforming network management from reactive adaptation to proactive optimization. The approach is enabled by fusing predictable mission trajectories with stable, large-scale radio environment models (e.g., radio maps). Specifically, we present a hierarchical framework that decomposes the predictive cross-layer resource allocation problem into three layers: strategic (routing), tactical (timing), and operational (power). This structure aligns decision-making timescales with the accuracy levels and ranges of available predictive information. We demonstrate that this foresight-driven framework achieves an order-of-magnitude reduction in cross-tier interference, laying the groundwork for robust and scalable low-altitude communication systems.
@@ I Introduction
The emergence of the low-altitude economy, an ecosystem of services conducted by unmanned aerial vehicles (UAVs) and electric vertical takeoff and landing (eVTOL) aircraft, is set to redefine urban and rural infrastructure [1, 2]. From automated logistics to real-time infrastructure inspection, these services depend on the seamless operation of extensive low-altitude networks. As these networks form a dynamic aerial layer above the terrestrial systems [3, 4], the critical role of a robust and intelligent wireless communication paradigm in low-altitude networks becomes paramount.
Figure 1: The operational concept of the low-altitude network and its implications for predictive communication. (a) Diverse low-altitude operations, such as infrastructure inspection and cargo delivery, are performed by aircraft following trajectories that are predetermined by their mission requirements. (b) For reasons of air traffic safety and operational management, these diverse, mission-driven flight plans are filed with a central control authority, which possesses a priori, network-wide knowledge of all aircraft movements. This centralized oversight of a predictable, time-varying network topology creates the foundational opportunity for predictive communication.
@@
  • We establish foundational principles of predictive communication in low-altitude networks, demonstrating the predictability of both trajectory and radio environment, and illustrating how these can be leveraged to forecast future channel states.
  • We analyze the hierarchical nature of predictive uncertainty, identifying a fundamental trade-off between predictive range and accuracy across central, local, and individual network echelons.
  • We propose a layered predictive communication framework, built upon the above hierarchical understanding. This framework decomposes the predictive cross-layer resource allocation problem into three layers: a strategic layer for network-level routing using coarse information, a tactical layer for group-level timing coordination, and an operational layer for precise link-level power allocation. A case study demonstrates the framework’s effectiveness in mitigating more than 10 dB of cross-tier interference.
@@ II-B Pillar 2: Radio Environment Predictability
The second pillar of our architecture is a predictable model of the radio environment that maps every transmitter-receiver coordinate pair to a statistical description of the large-scale channel. Formally, it is a static function from a 6D spatial input, the 3D coordinates for the transmitter and the receiver, to CSI statistics, including path-loss, shadowing, and etc. Two observations justify this pillar.
@@
  • Advanced radio-environment models now provide fine-grained, city-scale accuracy. State-of-the-art radio maps generated via ray tracing, measurement-driven interpolation, or high-fidelity digital twins can deliver metre-level resolution across the 3D workspace [12, 13, 14]
An aerial-assisted radio map is a powerful example of such a model, whose dynamic generation and update process is illustrated in Fig. 2. In this paradigm, aircraft act as mobile sensing agents, collecting geo-tagged channel measurements along their trajectories. This sparse dataset is collected and then transformed into a dense, multi-layered map at a central or edge repository using advanced radio-environment models that infer channel characteristics in unsampled locations. Crucially, this mechanism enables the continuous real-time update of the map, allowing it to adapt to environmental changes and reflect the most current state of the radio landscape.
@@ III Hierarchical Nature of Predictive Information
The predictive architecture outlined in the II provides an idealized glimpse of future connectivity; however, real-world forecasts are inherently probabilistic. Factors such as modeling errors, trajectory deviations, and feedback delays introduce uncertainty that grows with the prediction horizon. Moreover, the information available to different network entities, from a central controller to an individual aircraft, varies in scope, freshness, and fidelity, yielding a hierarchical information structure governed by a fundamental range-accuracy trade-off. This section identifies the principal sources of uncertainty, presents a three-tier hierarchy of predictive entities, and shows how their distinct sensing and processing capabilities lead to the trade-off summarized in Table I.
@@ III-A Fundamental Sources of Predictive Uncertainty
  • Trajectory deviations: Although missions are pre-planned, an aircraft’s realized path can drift from its filed trajectory because of wind gusts, temporary air-traffic directives, or minor navigation corrections. Even small lateral shifts are critical in the low altitude networks: a displacement of only a few meters may flip a link from line-of-sight (LOS) to NLOS, causing significant errors in large-scale channel prediction.
@@
Figure 5: Illustration of large-timescale proactive interference-aware path reservation. (a) The considered scenario, showing aircraft mission trajectories with directions (red arrows) within a ground layout that includes a source (square), a destination (triangle), and interference-sensitive nodes (circles). (b) The computed route and associated transmit power (colorbar) for a delay-tolerant task (delay tolerance is 20s). (c) The computed route and transmit power for a delay-sensitive task (delay tolerance is 2s). (d) shows the aggregate interference power of our proposed method versus two baselines under increasing network load.
@@ IV Layered Predictive Communication Framework
The analysis in the accuracy-range trade-off dictates a new design philosophy: one that embraces the hierarchical nature of predictive information as a structural blueprint for the optimization framework. This leads directly to our proposed solution: a layered predictive communication framework. The core principle of this framework is hierarchical task decomposition, as conceptually illustrated in Fig. 4. The predictive cross-layer objective of delivering data with minimal interference (Fig.  4(a)) is broken down into a cascade of distinct, tractable sub-problems. This is achieved by assigning different functional responsibilities to three layers, each operating at a different scale and leveraging the predictive information best suited to its task. This hierarchical decomposition is the key to creating a system that is both strategically robust and tactically agile.
@@ IV-A The Large-Scale Strategic Layer
The strategic layer, operating at the apex of the hierarchy, performs long-term, network-wide path reservation. This centralized function is critical, as only this layer possesses the global, long-horizon view necessary to optimize end-to-end routes, a task beyond the localized scope of lower echelons. While its predictive model is statistically coarse, it effectively establishes robust, high-level directives. The architecture is designed to dynamically refine this initial imprecision through the more agile tactical and reactive layers below. As illustrated in Fig. 4(b), the strategic layer translates global mission requirements into reserved paths using the global spatio-temporal channel graph, which serves as a foundational strategic directive for the entire network.
@@ IV-B The Middle-Scale Tactical Layer
Operating under the directives of the strategic layer, the tactical layer is responsible for coordinating multiple nodes within a localized context. This layer leverages more accurate, mid-term, local-area predictions to translate the high-level strategic plan into a refined set of concrete, coordinated actions. Its key function is to adapt the strategic plan to the dynamic realities of the local environment. A canonical example is illustrated in Fig. 4(c). The strategic layer may have issued a route directive of $2\to 6\to 7$. However, the nodes in the local cluster observe that the direct link $2\to 6$ is currently blocked or experiencing deep fading. Therefore, the tactical layer’s first action is to perform a local re-routing, adjusting the path to $2\to 5\to 7$ to bypass the obstruction. Having established a viable new path, it then proceeds to coordinate the transmission timing. Using their superior local predictive models, nodes 2 and 5 can identify and schedule their respective transmissions to coincide with periods of high channel gain. This dual capability of reactive re-routing and proactive re-timing allows the system to enhance efficiency and resilience in a way the coarse global model could not foresee.
@@ IV-C The Small-Scale Operational Layer
At the base of the hierarchy, the operational layer tackles short-term, link-level optimization. Each aircraft, equipped with on-board sensors, uses the most accurate measurement-anchored statistics to pre-compute a mapping from instantaneous channel states to physical-layer parameters. In Fig. 4(d), the node leverages high-fidelity predictions of its desired link and the air-to-ground interference link to derive an interference-aware power control rule ahead of transmission. By exploiting precise channel information, the operational layer enables low-latency, energy-efficient, and interference-aware resource allocation.
This layered decomposition is the key to creating a strategically robust and tactically agile system. The layers operate in a top-down cascade of goal-setting and refinement: the strategic layer’s global routing plan defines the high-level objectives and constraints for the tactical layer below. In turn, the tactical layer’s coordinated schedule provides a specific, timed mission for the operational layer. This structure enables the network to develop robust, long-term plans based on statistical foresight, while maintaining the flexibility to execute those plans with precision in response to real-world conditions. The practical efficacy of this hierarchical approach is validated in the subsequent case study.
V Case Study: Interference-Aware Predictive Communication
@@
This section validates our layered proactive framework through a comprehensive case study focused on predictive interference mitigation [15], with key results illustrated in Fig. 5. First, the large-scale layer demonstrates strategic adaptability, exploiting aircraft mobility for delay-tolerant tasks (Fig. 5(b)) while using multi-hop relays for delay-sensitive tasks (Fig. 5(c)). The middle-scale layer then performs tactical timing coordination based on channel dynamics, scheduling a high-power, rapid transmission for the degrading first two hops in Fig. 5(b) while using a low-power, longer transmission for the improving last hop. The small-scale layer ensures fine-grained interference control, strategically reducing transmit power when aircraft operate near sensitive ground nodes, a behavior confirmed by the colorbar. As validated in Fig. 5(d), this holistic, proactive strategy consistently achieves a significant reduction in aggregate interference power compared to traditional terrestrial (such as aggregate-based communication) and space (such as space-time-based communication) communication baselines, with this advantage becoming more pronounced as network load increases.
@@ VI Challenges and Opportunities
VI-A Dynamic Environmental Awareness: Real-Time and Semantic Model Updates
@@
VI-B Robust Decentralized Execution
While our framework provides a strategic plan, its execution can be made more robust against unforeseen events, such as node failures or unexpected blockages. A vital future direction is to enhance system resilience by moving towards decentralized execution models, likely using multi-agent reinforcement learning. The core research challenge is to develop decentralized policies that enable aircraft to follow the central plan and autonomously re-coordinate to find alternative solutions when disruptions occur, ensuring mission continuity. Techniques like knowledge distillation could be explored to accelerate this adaptive learning process.
VI-C Predictive Coverage
@@
VI-D Predictive integrated sensing and communication (ISAC)
@@ VII Conclusion
This paper pioneered a predictive communication paradigm for low-altitude networks, proposing a novel layered framework that translates hierarchical predictive information into robust strategies for routing, coordination, and policy design. To spur further innovation, several key future research directions were also identified to support and promote the continued development of this predictive framework.

### v2

@@ Abstract
The emergence of dense, mission-driven aerial networks supporting the low-altitude economy presents unique communication and security challenges, including extreme channel dynamics and severe cross-tier interference. Traditional reactive communication paradigms are ill-suited to these environments, as they fail to leverage the network’s inherent predictability. This paper introduces predictive communication, a novel paradigm transforming network management from reactive adaptation to proactive optimization. The approach is enabled by fusing predictable mission trajectories with stable, large-scale radio environment models (e.g., radio maps). Specifically, we present a hierarchical framework that decomposes the predictive cross-layer resource allocation problem into three layers: strategic (routing), tactical (timing), and operational (power). This structure aligns decision-making timescales with the accuracy levels and ranges of available predictive information. We demonstrate that this foresight-driven framework achieves an order-of-magnitude reduction in cross-tier interference and enables proactive security against threats such as jamming and spoofing, laying the groundwork for efficient, resilient, and secure low-altitude communication systems.
@@ I Introduction
The emergence of the low-altitude economy, an ecosystem of services conducted by unmanned aerial vehicles (UAVs) and electric vertical takeoff and landing (eVTOL) aircraft, is set to redefine urban and rural infrastructure [1, 2]. From automated logistics to real-time infrastructure inspection, these services depend on the seamless operation of extensive low-altitude networks. As these networks form a dynamic aerial layer above the terrestrial systems [3, 4], a communication paradigm that ensures efficient, reliable, and secure operation becomes paramount.
Figure 1: The operational concept of the low-altitude network and its implications for predictive communication. (a) Diverse low-altitude operations, such as infrastructure inspection and cargo delivery, are performed by aircraft following predetermined trajectories that are tailored to their mission requirements. (b) For reasons of air traffic safety and operational management, these diverse, mission-driven flight plans are filed with a central control authority, which possesses a priori, network-wide knowledge of all aircraft movements. This centralized oversight of a predictable, time-varying network topology creates the foundational opportunity for predictive communication.
@@
  • We formalize the concept of predictive communication for low-altitude networks, establishing its feasibility upon two pillars: a priori mission trajectories and large-scale radio environment models. We also analyze the hierarchical nature of this predictive information, revealing a fundamental trade-off between predictive range and accuracy that governs system design.
  • We propose a novel layered predictive communication framework that is structurally matched to this information hierarchy. The framework decomposes the complex network management problem into three distinct layers: strategic long-term routing, tactical mid-term coordination, and operational short-term policy optimization, enabling a robust and efficient use of imperfect foresight.
  • We demonstrate our framework’s dual utility. Its effectiveness is quantitatively validated in a case study on interference mitigation, while its versatility is conceptually demonstrated by outlining how the same architecture, through tailored objectives and constraints, can proactively address security threats like jamming, spoofing, and interception.
The rest of this article is organized as follows. We introduce the predictive information and its hierarchical nature. Then, the proposed framework and security-oriented extension are presented. Finally, we provide some future directions and conclude the article.
@@ II-B Pillar 2: Radio Environment Predictability
The second pillar of our architecture is a predictable model of the radio environment that maps every transmitter-receiver coordinate pair to a statistical description of the large-scale channel. Formally, it is a static function from a 6D spatial input, the 3D coordinates for the transmitter and the receiver, to CSI statistics, including path loss, shadowing, etc. Two observations justify this pillar.
@@
  • Advanced radio-environment models now provide fine-grained, city-scale accuracy. State-of-the-art radio maps generated via ray tracing, measurement-driven interpolation, or high-fidelity digital twins can deliver metre-level resolution across the 3D workspace [12, 13, 14].
An aerial-assisted radio map is a powerful example of such a model, and its generation and update processes are illustrated in Fig. 2. In this paradigm, aircraft act as mobile sensing agents, collecting geo-tagged channel measurements along their trajectories. This sparse dataset is collected and then transformed into a dense, multi-layered map at a central or edge repository using advanced radio-environment models that infer channel characteristics in unsampled locations. Crucially, this mechanism enables the continuous real-time update of the map, allowing it to adapt to environmental changes and reflect the most current state of the radio landscape.
@@ III Hierarchical Nature of Predictive Information
The predictive architecture outlined in Section II provides an idealized glimpse of future connectivity; however, real-world forecasts are inherently probabilistic. Factors such as modeling errors, trajectory deviations, and feedback delays introduce uncertainty that grows with the prediction horizon. Moreover, the information available to different network entities, from a central controller to an individual aircraft, varies in scope, freshness, and fidelity, yielding a hierarchical information structure governed by a fundamental range-accuracy trade-off. This section identifies the principal sources of uncertainty, presents a three-tier hierarchy of predictive entities, and shows how their distinct sensing and processing capabilities lead to the trade-off summarized in Table I.
@@ III-A Fundamental Sources of Predictive Uncertainty
  • Trajectory deviations: Although missions are pre-planned, an aircraft’s realized path can drift from its filed trajectory due to wind gusts, temporary air-traffic directives, minor navigation corrections, or proactive security maneuvers, such as detouring around a newly identified high-risk area. Even small lateral shifts are critical in the low altitude networks: a displacement of only a few meters may flip a link from line-of-sight (LOS) to non-line-of-sight (NLOS), causing significant errors in large-scale channel prediction.
@@
Figure 5: Interference-aware predictive communications. (a) Case study. (a.1) Considered scenario with predictive aircraft mission trajectories, where black dots denote aircraft positions and red arrows indicate flight directions, over a ground layout comprising a source (square), a destination (triangle), and interference-sensitive nodes (circles). (a.2) Planned route and associated transmit power (color bar) for a delay-tolerant task (delay tolerance 20 s). (a.3) Planned route and transmit power for a delay-sensitive task (delay tolerance 2 s). (b) Aggregate interference power of the proposed method versus two baselines under increasing network load (number of commodities).
@@ IV Layered Predictive Communication Framework
The core principle of this framework is hierarchical task decomposition, as conceptually illustrated in Fig. 4. The predictive cross-layer objective of delivering data with minimal interference (Fig. 4(a)) is broken down into a cascade of distinct, tractable sub-problems. This is achieved by assigning different functional responsibilities to three layers, each operating at a different scale and leveraging the predictive information best suited to its task. Within this hierarchy, channel uncertainty caused by trajectory deviations, environmental modeling errors, and small-scale fading is captured through a statistical characterization of the problem, e.g., expectation- or outage-based objectives and constraints.
@@ IV-A The Large-Scale Strategic Layer
The strategic layer performs long-term, network-wide path reservation. Operating at the apex of the hierarchy, it leverages a global, long-horizon predictive view to optimize end-to-end routes, a task impossible for lower echelons with only localized information. While its predictive model is statistically coarse, it is uniquely capable of establishing robust, high-level directives for the entire network. For example, as illustrated in Fig. 4(b), the strategic layer translates global mission requirements into reserved paths using the global spatio-temporal channel graph. By incorporating threat intelligence into the planning process, the strategic layer can steer entire data flows around known jamming zones or regions with high interception risk, explicitly trading off path length for improved security.
@@ IV-B The Middle-Scale Tactical Layer
The tactical layer adapts the strategic plan to dynamic local realities using more accurate, mid-term predictions. Its key functions are multi-node coordination, reactive re-routing, and proactive re-timing. A canonical example is illustrated in Fig. 4(c). The strategic layer may have issued a route directive of $2\to 6\to 7$. However, the nodes in the local cluster observe that the direct link $2\to 6$ is physically blocked by a building or virtually blocked by a suspected jamming attack or a newly imposed "no-fly" zone. Therefore, the tactical layer’s first action is to perform a local re-routing, adjusting the path to $2\to 5\to 7$ to bypass the obstruction. Having established a viable new path, it then proceeds to coordinate the transmission timing. Using their superior local predictive models, nodes 2 and 5 can identify and schedule their respective transmissions to coincide with periods of high channel gain and low security exposure. This dual capability of reactive re-routing and proactive re-timing allows the system to enhance efficiency and resilience in a way the coarse global model could not foresee.
@@ IV-C The Small-Scale Operational Layer
At the base of the hierarchy, the operational layer tackles short-term, link-level optimization. Using the most accurate, measurement-anchored statistics, each aircraft pre-computes an optimal transmission policy. In Fig. 4(d), the node leverages high-fidelity predictions of its desired link and the air-to-ground interference link to derive an interference-aware power control rule ahead of transmission. This same predictive mechanism is vital for enhancing confidentiality. By also predicting the channel to a potential eavesdropper, the aircraft can design a low probability of intercept power policy, dynamically adjusting its transmit power to meet its own communication needs while remaining below the eavesdropper’s detection threshold. This enables energy-efficient, and secure resource allocation.
This layered decomposition is the key to a strategically robust yet tactically agile system. The strategic layer uses long-term predictive information to set global routing goals and constraints. In contrast, the tactical and operational layers refine these plans with progressively shorter prediction-execution gaps using fresher channel observations. In this way, long-horizon decisions are continuously corrected at lower layers, so the network remains robust to trajectory and environmental uncertainties while reacting quickly to real-world conditions.
V Case Study and Security Extensions
@@
This section validates our layered predictive framework through a case study on interference mitigation [15], with key results illustrated in Fig. 5. First, the large-scale layer demonstrates strategic adaptability, exploiting aircraft mobility for delay-tolerant tasks (Fig. 5(a.2)) while using multi-hop relays for delay-sensitive tasks (Fig. 5(a.3)). The middle-scale layer then performs tactical timing coordination based on channel dynamics, scheduling a high-power, rapid transmission for the degrading first two hops in Fig. 5(a.2) while using a low-power, longer transmission for the improving last hop. The small-scale layer ensures fine-grained interference control, strategically reducing transmit power when aircraft operate near sensitive ground nodes. Compared with the classical aggregate-based baseline (routing on a static graph with locally coordinated handovers) and the space-time-based baseline (routing on a space-time graph with fixed, equally spaced handover times), the proposed layered strategy consistently achieves a larger reduction in aggregate interference power, as validated in Fig. 5(b), with this advantage becoming more pronounced as network load increases. Note that goal-oriented network formulations reduce the computational burden and keep centralized implementation tractable.
Although this case study focuses on interference-aware communications, the same layered predictive framework is also well-suited for security-oriented designs in low-altitude networks. By appropriately augmenting the optimization objectives and constraints with security-related metrics, predictive routing, handover timing, and power control can be tailored to minimize exposure to suspected jamming or high-risk regions in space and time; radio-map-assisted channel prediction can serve as a prior for anomaly detection in positioning and control signals (e.g., GPS spoofing or link hijacking); and the hierarchical timing structure naturally supports geofencing- and time-aware scheduling policies that avoid transmitting sensitive data over vulnerable links or areas, without altering the core architecture of the framework.
@@ VI Challenges and Opportunities
VI-A Dynamic Environmental Awareness
@@
VI-B Predictive Coverage
@@
VI-C Predictive Integrated Sensing and Communication (ISAC)
@@
VI-D Predictive Security and Resilience
While this paper introduces the potential for predictive security, it opens a rich field for future work. Research is needed to develop sophisticated threat models that can be integrated into the predictive objectives, moving beyond simple high-risk zones to probabilistic models of attacker behavior. Furthermore, leveraging the framework for distributed, multi-agent security games, where aircraft learn to cooperatively detect and isolate threats in a decentralized manner, represents a vital direction for enhancing network resilience.
@@ VII Conclusion
This paper proposes a predictive communication framework for low-altitude networks, which leverages hierarchical predictive information to derive robust strategies for routing, coordination, and policy design. To spur further innovation, we outline several key future research directions that will support and promote the continued development of this framework.